Consider a 3-factor Arbitrage Pricing Theory (APT) model. Assuming a risk-free rate of 4%, calculate the...

  1. Consider a 3-factor Arbitrage Pricing Theory (APT) model. Assuming a risk-free rate of 4%, calculate the expected return of this stock.                                                                                           

Factor

Risk Premium

Sensitivity to each factor

Change in GDP

5%

1

Change in interest rate

1%

0.5

Inflation ratio

2.5%

0.2

(4 marks)

  1. Consider the following portfolio composed of 3 stocks (A, B, C):

Stock

Quantity

Price (£)

Beta

A

500

1.5

0.8

B

520

1.7

0.97

C

610

1.1

1.04

What is the beta of this portfolio?                                                             (4 marks)

  1. Explain the security market line in the context of the CAPM and examine how this can be used to identify whether a security is under-priced or over-priced.                                                                                                      (5 marks)

  1. Critically examine the empirical evidence in relation to the CAPM.                                                                                                                           (8 marks)
Answer #1

Solution :-

(A)

Expected Return of Stock Under APT =

Risk Free Rate + ( Risk Premium of Various Factors*Sensitivity of Factor )

= 0.04 + (0.05*1 + 0.01*0.5 + 0.025*0.2)

= 0.04 + 0.06 = 10%

(B)

(c) Security market line is simple we say a graphical representation of CAPM of different securities on the basis of systematic risk.

An overpriced security will have a lower return than SML and Underpriced has a more return than SML.

Brother this answer is too long so ask other last part as seperate one because i dont have time now and also it is policy of HomeworkLib to answer one question at once